# Perfect Matching in Product Graphs and in their Random Subgraphs

### v2

@@ Abstract
For $t\in\mathbb{N}$ and every $i\in[t]$, let $H_{i}$ be a $d_{i}$-regular connected graph, with $1<|V(H_{i})|\leq C$ for some integer $C\geq 2$. Let $G=\square_{i=1}^{t}H_{i}$ be the Cartesian product of $H_{1},\ldots,H_{t}$. We show that if $t\geq 5C\log_{2}C$ then $G$ contains a (nearly-)perfect matching.
Then, considering the random graph process on $G$, we generalise the result of Bollobás on the binary hypercube $Q^{t}$, showing that with high probability, the hitting times for minimum degree one, connectivity, and the existence of a (nearly-)perfect matching in the $G$-random-process are the same. We develop several tools which may be of independent interest in a more general setting of the typical existence of a perfect matching under percolation.
@@ 1.1 Background and main results
$\left\{uv\colon\begin{array}[]{l}\text{there is some }i\in[t]\text{ such that }u_{j}=v_{j}\\ \text{for all }i\neq j\text{ and }\{u_{i},v_{i}\}\in E\left(H_{i}\right)\end{array}\right\}.$
@@
In this paper, we study perfect matchings in product graphs. One can convince oneself that if one of the base graphs has a perfect matching, then the product graph $G$ has a perfect matching as well. Indeed, Kotzig [19] showed something even stronger: let $G=\square_{i=1}^{t}H_{i}$ be such that each $H_{i}$ is regular, then if at least one of the base graphs has a $1$-factorisation (recall that a $1$-factorisation is a decomposition of the edges of the graph into disjoint perfect matchings), or at least two base graphs have perfect matchings, then $G$ has a $1$-factorisation; and furthermore these sufficient conditions are not necessary. However, when none of the base graphs have a perfect matching, it is not clear whether their product graph will have a perfect matching. Indeed, not every connected regular graph has a perfect matching (or nearly-perfect matching, that is, a matching covering all but one vertex when the order of the graph is odd). However, note that given two graphs $H_{1}$ and $H_{2}$, the proportion of vertices covered by a largest matching is at least as large as this proportion in any of the base graphs. Thus, taking the product of many graphs, one can hope to improve this proportion. Indeed, our first result gives a simple sufficient condition for the existence of a perfect matching in product graphs in this spirit.
@@
Let $C>1$ be a constant, and let $t\geq 5C\log_{2}C$ be an integer. For every $i\in[t]$, let $H_{i}$ be a $d_{i}$-regular connected graph with $1<|V(H_{i})|\leq C$. Let $G=\square_{i=1}^{t}H_{i}$ and suppose that $|V(G)|$ is even. Then $G$ has a perfect matching.
Observe that since each $H_{i}$ is a $d_{i}$-regular connected graph, $G=\square_{i=1}^{t}H_{i}$ is a connected $d$-regular graph with $d=\sum_{i=1}^{t}d_{i}$, and in particular $d=\Theta_{C}(\log|V(G)|)$. Furthermore, we note that we have not tried to optimise the constants in the above. Theorem 1 shows that if the dimension of the product graph is sufficiently large, with respect to the maximum size of the base graphs, then the product graph has a perfect matching — regardless of whether the base graphs themselves contain any perfect matchings. Furthermore, from the same proof it follows that if $G$ has an odd number of vertices, then it has a nearly-perfect matching.
As mentioned above, a well-studied product graph is the binary hypercube, $Q^{t}$, in particular in terms of bond percolation on it. Given a graph $G$ and probability $p$, we form the percolated random subgraph $G_{p}\subseteq G$ by including every edge of $G$ independently with probability $p$ (note that $G(n,p)$ is then $(K_{n})_{p}$). The study of $Q^{t}_{p}$ has been initiated by Sapoženko [22] and by Burtin [8], who showed that the sharp threshold for connectivity is $p^{*}=\frac{1}{2}$: when $p<\frac{1}{2}$, with high probability (whp), that is, with probability tending to one as $t$ tends to infinity, $Q^{t}_{p}$ is disconnected, whereas for $p>\frac{1}{2}$, whp $Q^{t}_{p}$ is connected. Erdős and Spencer [14] conjectured that $Q^{t}_{p}$ undergoes a phase transition with respect to its component structure, that is, the typical emergence of a giant component (a connected component containing a linear fraction of the vertices) around $p=\frac{1}{t}$, similar to that of $G(n,p)$ around $p=\frac{1}{n}$. This conjecture was confirmed by Ajtai, Komlós, and Szemerédi [1], with subsequent work by Bollobás, Kohayakawa, and Łuczak [6]. We refer the reader to [20] for a modern short proof of this result.
@@
Returning to the dense regime, that is, when $p$ is constant, connectivity and the existence of a perfect matching in $Q^{t}_{p}$ have been studied in detail. In particular, Bollobás [4] obtained a hitting time result for the random graph process on $Q^{t}$. Given a graph $\Gamma$, the random graph process on $\Gamma$ is defined as a random sequence of nested graphs $\Gamma(0)\subseteq\ldots\subseteq\Gamma(|E(\Gamma)|)$ together with an ordering $\sigma$ on $E(\Gamma)$, chosen uniformly at random from among all $|E(\Gamma)|!$ such orderings. We set $\Gamma(0)$ to be the empty graph on $V(\Gamma)$. Given $\Gamma(i)$, with $0\leq i<|E(\Gamma)|$, we form $\Gamma(i+1)$ by adding the $(i+1)$-th edge, according to the ordering $\sigma$, to $\Gamma(i)$. The hitting time of a monotone increasing, non-empty graph property $\mathcal{P}$, is a random variable equal to the index $\tau$ for which $\Gamma(\tau)\in\mathcal{P}$, but $\Gamma(\tau-1)\notin\mathcal{P}$. Note that having minimum degree one, connectivity, and the existence of a perfect matching are all monotone increasing properties. Furthermore, observe that for a graph $\Gamma$ to be connected or to contain a perfect matching, the minimum degree of $\Gamma$ has to be at least one. A classical result of Erdős and Rényi [13], and of Bollobás and Thomason [7], is that in the random process on $K_{n}$, whp the hitting time for minimum degree one, connectivity, and the existence of a perfect matching are the same. In 1990, Bollobás [4] showed that the same phenomenon holds in the random graph process on $Q^{t}$.
@@
Subsequent work by Joos determined the threshold for connectivity for Cartesian powers of graphs [17, 18], that is, $\square_{i=1}^{t}H_{i}$ where all the base graphs $H_{i}$ are the same.
@@
In fact, we prove the following description of $G_{p}$, from which standard results (see, for example, [5, Theorem 7.4] and [4, Lemma 2]), allow one to derive Theorem 2.
@@
Let $C>1$ be an integer. For every $i\in[t]$, let $H_{i}$ be a $d_{i}$-regular connected graph with $1<|V(H_{i})|\leq C$. Let $G=\square_{i=1}^{t}H_{i}$, and suppose that $n\coloneqq|V(G)|$ is even. Let $d\coloneqq d(G)=\sum_{i=1}^{t}d_{i}$. Let $\epsilon\geq 0$ be a sufficiently small constant, and let $p$ be such that $(1-p)^{d}=n^{-(1-\epsilon)}$. Then, whp, the following holds in $G_{p}$.
@@
A few comments are in place. Note that the threshold probability $p^{*}$, at which the expected number of isolated vertices is zero, is $p^{*}=1-\left(\frac{1}{n}\right)^{1/d}$. Noting that $t\leq d\leq(C-1)t$, we have that $d=\Theta_{C-1}(\log n)$. Thus, $\left(\frac{1}{n}\right)^{1/d}$ is bounded away from zero, and bounded away from one, and in turn, $p^{*}$ is bounded away both from zero and from one. Further, if $p$ is such that $(1-p)^{d}=n^{-(1-\epsilon)}$ for small enough constant $\epsilon\geq 0$, then $p$ is bounded away both from zero and from one.
Furthermore, let us remark that Theorem 3a implies that whp adding any edge to $G_{p}$, which is not contained in the giant component, must connect an isolated vertex to the giant component, and indeed this suffices to show that the hitting times of minimum degree one and connectivity are whp the same. Theorem 3b then shows that whp the only obstacle to a perfect matching in $G_{p}$ itself is the existence of isolated vertices. Moreover, the proof of Theorem 3a does not rely on the product structure of the graph, and only uses the assumptions that the graph is $d$-regular, has optimal edge-expansion for sets of size polynomial in $d$, and has some mild edge-expansion for larger sets. Only in the proof of Theorem 3b we use the product structure explicitly.
Let us briefly remark that since the hypercube is bipartite, in order to find a perfect matching it suffices to apply Hall’s theorem. In our setting, however, the graph $G$ is not necessarily bipartite, which means that in order to show the typical existence of a perfect matching, one needs to consider the Tutte-Berge formula. This requires a much more delicate treatment, which is detailed in the proof outline in the subsequent section, and also requires us to develop tools which are valid for more general graphs (see Lemmas 3.2 and 3.7) and could be of independent interest when treating perfect matchings in the setting of percolation.
@@ 1.2 Proofs outline
Throughout the paper we let $C>1$ be an integer, $(H_{i})_{i=1}^{t}$ be a sequence of $d_{i}$-regular and connected graphs with $1<|V(H_{i})|\leq C$, and let $G=\square_{i=1}^{t}H_{i}$ be their product. We let $G=(V,E)$. We call $t$ the dimension of $G$, and given $u=(u_{1},\ldots,u_{t})\in V$ we call $u_{i}$ the $i$-th coordinate of $u$. Furthermore, we denote by $n:=|V|$ the order of $G$, and by $d\coloneqq d(G)$ the degree of $G$, noting once again that $d=\sum_{i=1}^{t}d_{i}=\Theta_{C}(\log n)$.
@@
Finally, throughout the paper, we let $\epsilon\geq 0$ be a sufficiently small constant, and let $p$ be such that $(1-p)^{d}=n^{-(1-\epsilon)}$, recalling that this implies that $p=\Theta_{C}(1)$.
@@
$e(S,S^{C})\geq\frac{|S|\log_{C}\left(\frac{n}{|S|}\right)}{C-1}\text{ and }$
$e(S,S^{C})\geq|S|(d-(C-1)\log_{2}|S|).$
@@
The proof follows from a double-exposure argument, similar in spirit to the classical argument of Ajtai, Komlós, and Szemerédi [1]. Let $p_{2}\coloneqq\frac{1}{d^{2}}$ and let $p_{1}$ be such that $(1-p_{1})(1-p_{2})=1-p$. Note that $G_{p_{1}}\cup G_{p_{2}}\sim G_{p}$, and $p_{1}\approx p-\frac{1}{d^{2}}$. We first show, using a first-moment calculation, that whp every two vertices $u,v$ which are isolated in $G_{p_{1}}$ are at distance at least two in $G$ (Lemma 2.1). Then, utilising Theorem 1.2 and Lemma 1.3, we show that whp there are no components in $G_{p_{1}}$ whose order is in $\left[2,d^{20}\right]$ (Lemma 2.2). We then turn to show that typically components whose order is at least $d^{20}$ merge after sprinkling with $p_{2}$ — this is fairly standard, and follows the same approach as in [1]. Noting that any vertex which was an isolated vertex in $G_{p_{1}}$, is whp either an isolated vertex after sprinkling with $p_{2}$ or merged into a component which was of size at least $d^{20}$ in $G_{p_{1}}$, we obtain that whp $G_{p}$ has a unique large connected component, and all the other vertices are isolated vertices. Finally, from Markov’s inequality it follows that the total volume of isolated vertices in $G_{p}$ is at most $n^{2\epsilon}=o(n)$.
@@
Recall that we seek to show that the giant component of $G_{p}$ has a (nearly-)perfect matching. We know from Theorem 3a that whp every vertex of $G$ which is not in the giant component is an isolated vertex in $G_{p}$. Thus, it suffices to show that whp $G_{p}$ has a perfect matching if there are typically no isolated vertices in $G_{p}$. To that end, we will utilise the Tutte-Berge formula [2], which implies that the number of unmatched vertices in a maximum matching in a graph $\Gamma$ is equal to
@@
We thus define Tutte-like obstructions — we say that $U\subseteq V$ is an obstruction (for $G_{p}$) if $|U|\geq 1$, and the number of components of size different than two in $G_{p}[V\setminus U]$ is at least $|U|+1$. We say that the size of the obstruction is $|U|$. Note that if $G_{p}$ has no such obstructions, then the only possible obstructions to a perfect matching are isolated vertices, and whp the giant component has a perfect matching. Furthermore, observe that this definition does not capture all the properties coming from the Tutte-Berge formula – indeed, we treat components with an even number of vertices, except two, the same as components with an odd number of vertices – but this crude, yet much-simplified, outlook will suffice for our needs.
We denote the components in $G_{p}[V\setminus U]$ of size different from two by $K_{1},\ldots,K_{\ell}$, where $\ell\geq|U|+1$ if $U$ is an obstruction, and let $K=\bigcup_{i=1}^{\ell}K_{i}$ and $k_{i}\coloneqq|K_{i}|$ for every $i\in[\ell]$. We denote the components of size two in $G_{p}[V\setminus U]$ by $W_{1},\ldots,W_{w/2}$, such that their total volume is $w$.
@@
We say that $U$ is a minimal obstruction, if it is an obstruction with the smallest size of $U$. Note that if a graph has an obstruction, then it must also have a minimal obstruction. We will use the convention that $u\coloneqq|U|$ and $k\coloneqq|K|$. Note that $u+k+w=n$.
@@
We will require a finer description of the components in $K$. Let us partition the vertices in $K$ into three sets, according to the size of the component in $G_{p}[V\setminus U]$ they belong to. Let $\ell_{1}$ be the number of components of size one in $G_{p}[V\setminus U]$, let $\ell_{2}$ be the number of components whose size is in $[3,d^{2}]$ in $G_{p}[V\setminus U]$, and let $\ell_{3}=\ell-\ell_{1}-\ell_{2}$ be the number of components in $G_{p}[V\setminus U]$ that contain more than $d^{2}$ vertices. We further denote the set of vertices in components of $G_{p}[V\setminus U]$ whose size is one by $V_{1}$, those whose size is in $[3,d^{2}]$ by $S$, and those whose size is larger than $d^{2}$ by $B$. We write $s:=|S|$ and $b:=|B|$, and note that $\ell_{1}=|V_{1}|$.
@@
Figure 1: Illustration of an obstruction, with the sets $U$, $V_{1}$, $W$, $S$ and $B$. Note that the only edges in $G_{p}$, which are not induced by $U$ or by components $K_{i}$ in $K$ or $W_{i}$ in $W$, are in $E(U,V_{1}\cup W\cup S\cup B)$.
@@
Note that an obstruction is uniquely determined by the choice of $U$, and that $U$ is not necessarily connected, and thus the number of choices could be $\binom{n}{u}$. As is often the case, a key ingredient will be to efficiently enumerate the number of possible obstructions, and $\binom{n}{u}$ will often be an inefficient bound. We thus prove two claims, which could be of independent interest in arguing about perfect matchings under percolation for general graphs. First, in Lemma 3.2, we show that a minimal obstruction of size $u$ is determined by the choice of $W\cup S\cup B$. We stress that this result holds for any graph $G$, without requirements on its degree or structure. As these sets contain connected components, by Lemma 1.3 there are at most (roughly) $n^{w/2+\ell_{2}+\ell_{3}}(ed)^{s+w+b}$ choices for a minimal obstruction of size $u$ (see Lemma 3.3 for a precise statement). Since $d=\Theta_{C}(\log n)$, for ‘large’ $u$ this will be a much more efficient bound than $\binom{n}{u}$. As it turns out, for certain types of obstructions, neither of these bounds will be efficient enough, and we will require a different approach. Utilising the isoperimetric properties of $G$ (Theorem 1.2), we show in Lemma 3.7 that whp if $U$ is an obstruction of order polynomial in $n$ (that is, $n^{a}$ for some constant $10<a<$ 1), then there is a set $M$ such that $U\subseteq M$, $|M|\leq 10u$, and there are relatively few connected components in $M$ — that is, $U$ must reside in a relatively well-connected set of order proportional to $|U|$. This allows us to more efficiently bound the number of choices of such obstructions, using Lemma 1.3.

### v3

@@ Abstract
For $t\in\mathbb{N}$ and every $i\in[t]$, let $H_{i}$ be a $d_{i}$-regular connected graph, with $1<|V(H_{i})|\leq C$ for some integer $C\geq 2$. Let $G=\square_{i=1}^{t}H_{i}$ be the Cartesian product of $H_{1},\ldots,H_{t}$. We show that if $t\geq 5C$ then $G$ contains a (nearly-)perfect matching.
Then, considering the random graph process on $G$, we generalise the result of Bollobás on the binary hypercube $Q^{t}$, showing that with high probability, the hitting times for minimum degree one, connectivity, and the existence of a (nearly-)perfect matching in the random graph process on $G$ are the same. As a byproduct, we develop several tools which may be of independent interest in a more general setting when one seeks to establish the typical existence of a perfect matching under percolation.
@@ 1.1 Background and main results
[…]
@@
In this paper, we study perfect matchings in product graphs. One can convince oneself that if one of the base graphs has a perfect matching, then the product graph $G$ has a perfect matching as well. Indeed, Kotzig [19] showed something even stronger: let $G=\square_{i=1}^{t}H_{i}$ be such that each $H_{i}$ is regular, then if at least one of the base graphs has a $1$-factorisation (recall that a $1$-factorisation is a decomposition of the edges of the graph into disjoint perfect matchings), or at least two base graphs have perfect matchings, then $G$ has a $1$-factorisation; he furthermore demonstrated that while these conditions are sufficient, they are not necessary. Indeed, when none of the base graphs have a perfect matching, it is not clear whether their product graph will have a perfect matching (or nearly-perfect matching, that is, a matching covering all but one vertex when the order of the graph is odd). However, note that given two graphs $H_{1}$ and $H_{2}$, the proportion of vertices covered by a largest matching in $H_{1}\square H_{2}$ is at least as large as this proportion in any of the base graphs. Thus, taking the product of many graphs, one can hope to improve this proportion. Indeed, our first result gives a simple sufficient condition for the existence of a perfect matching in product graphs in this spirit.
@@
Let $C>1$ be a constant, and let $t\geq 5C$ be an integer. For every $i\in[t]$, let $H_{i}$ be a $d_{i}$-regular connected graph with $1<|V(H_{i})|\leq C$. Let $G=\square_{i=1}^{t}H_{i}$ and suppose that $|V(G)|$ is even. Then $G$ has a perfect matching.
Observe that since each $H_{i}$ is a $d_{i}$-regular connected graph, $G=\square_{i=1}^{t}H_{i}$ is a connected $d$-regular graph with $d=\sum_{i=1}^{t}d_{i}$, and in particular $d=\Theta_{C}(\ln|V(G)|)$. Theorem 1 shows that if the dimension of the product graph is sufficiently large with respect to the maximum size of the base graphs, then the product graph has a perfect matching — regardless of whether the base graphs themselves contain any perfect matchings. Furthermore, we note that the same proof gives that if $|V(G)|$ is odd, then $G$ has a nearly-perfect matching. Finally, the following remark shows that some regularity assumption on the base graphs is necessary.
Remark 1.1.
Note that Theorem 1 is not true if we drop the assumption that the base graphs are regular. Indeed, let $H=K_{1,s}$ be a star and consider the product of $t$ stars $G=\square_{i=1}^{t}H$. Since $K_{1,s}$ is bipartite, $G$ is bipartite for any $t$. However, for any $t\geq 1$, this bipartition is never balanced. In particular, the graph $G$ does not contain a (nearly-)perfect matching. More precisely, denote by $\mathcal{O},\mathcal{E}$ the bipartition sides. Then
$|\mathcal{O}(G)|=\sum_{\begin{subarray}{c}k=0\\ k\text{ even}\end{subarray}}^{t}\binom{t}{k}s^{k}\qquad\text{and}\qquad|\mathcal{E}(G)|=\sum_{\begin{subarray}{c}k=1\\ k\text{ odd}\end{subarray}}^{t}\binom{t}{k}s^{k}.$
The number of vertices in $G$ is $(s+1)^{t}$ and the difference between the bipartition sides is $|\mathcal{O}(G)|-|\mathcal{E}(G)|=(1-s)^{t}$. This is strictly larger than one for any $t\geq 1$ if $s\geq 3$.
As mentioned above, a well-studied product graph is the binary hypercube, $Q^{t}$, in particular in terms of bond percolation on it. Given a graph $G$ and probability $p$, we form the percolated random subgraph $G_{p}\subseteq G$ by including every edge of $G$ independently with probability $p$ (note that $G(n,p)$ is then $(K_{n})_{p}$). The study of $Q^{t}_{p}$ was initiated by Sapoženko [22] and by Burtin [7], who showed that the (sharp) threshold for connectivity is $p^{*}=\frac{1}{2}$: when $p<\frac{1}{2}$ is a constant, with high probability (whp), that is, with probability tending to one as $t$ tends to infinity, $Q^{t}_{p}$ is disconnected, whereas for $p>\frac{1}{2}$ constant, whp $Q^{t}_{p}$ is connected. Erdős and Spencer [14] conjectured that $Q^{t}_{p}$ undergoes a phase transition with respect to its component structure, that is, the typical emergence of a giant component (a connected component containing a linear fraction of the vertices) around $p=\frac{1}{t}$, similar to that of $G(n,p)$ around $p=\frac{1}{n}$. This conjecture was confirmed by Ajtai, Komlós, and Szemerédi [1], with subsequent work by Bollobás, Kohayakawa, and Łuczak [5]. We refer the reader to [20] for a modern short proof of this result.
@@
Returning to the dense regime, that is, when $p$ is constant, connectivity and the existence of a perfect matching in $Q^{t}_{p}$ have been studied in detail. In particular, Bollobás [4] obtained a hitting time result for the random graph process on $Q^{t}$. Given a graph $\Gamma$, the random graph process on $\Gamma$ is defined as a random sequence of nested graphs $\Gamma(0)\subseteq\ldots\subseteq\Gamma(|E(\Gamma)|)$ together with an ordering $\sigma$ on $E(\Gamma)$, chosen uniformly at random from among all $|E(\Gamma)|!$ such orderings. We set $\Gamma(0)$ to be the empty graph on $V(\Gamma)$. Given $\Gamma(i)$, with $0\leq i<|E(\Gamma)|$, we form $\Gamma(i+1)$ by adding the $(i+1)$-th edge, according to the ordering $\sigma$, to $\Gamma(i)$. The hitting time of a monotone increasing, non-empty graph property $\mathcal{P}$, is the random variable equal to the index $\tau$ for which $\Gamma(\tau)\in\mathcal{P}$, but $\Gamma(\tau-1)\notin\mathcal{P}$. Note that having minimum degree one, connectivity, and the existence of a perfect matching are all monotone increasing properties. Furthermore, observe that for a graph $\Gamma$ to be connected or to contain a perfect matching, the minimum degree of $\Gamma$ has to be at least one. A classical result of Erdős and Rényi [13], and of Bollobás and Thomason [6], is that in the random process on $K_{n}$, whp the hitting time for minimum degree one, connectivity, and the existence of a perfect matching are the same. In 1990, Bollobás [4] showed that the same phenomenon holds in the random graph process on $Q^{t}$.
@@
Subsequent work by Joos determined the threshold for connectivity for Cartesian powers of graphs [17, 18], that is, for $G=\square_{i=1}^{t}H$ where all the base graphs $H_{i}=H$ are the same.
@@
In fact, we prove the following description of $G_{p}$, from which we derive Theorem 2.
@@
Let $C>1$ be an integer. For every $i\in[t]$, let $H_{i}$ be a $d_{i}$-regular connected graph with $1<|V(H_{i})|\leq C$. Let $G=\square_{i=1}^{t}H_{i}$, and suppose that $n\coloneqq|V(G)|$ is even. Let $d\coloneqq d(G)=\sum_{i=1}^{t}d_{i}$. Let $\omega=\omega(d)$ be a function tending arbitrarily slowly to infinity, and let $p$ be such that $(1-p)^{d}=\omega/n$. Then, whp, the following holds in $G_{p}$.
@@
A few comments are in place. Note that the threshold probability $p^{*}$, at which the expected number of isolated vertices is zero, is $p^{*}=1-\left(\frac{1}{n}\right)^{1/d}$. Noting that $t\leq d\leq(C-1)t$, we have that $d=\Theta_{C-1}(\ln n)$. Thus, $\left(\frac{1}{n}\right)^{1/d}$ is bounded away from zero, and bounded away from one, and in turn, $p^{*}$ is bounded away both from zero and from one. Further, if $p$ is such that $(1-p)^{d}=\omega/n$ for $\omega$ tending to infinity arbitrarily slowly, then $p$ is bounded away both from zero and from one.
Note, that for our choice of $p$ such that $(1-p)^{d}=\omega/n$, whp, there are still isolated vertices. Furthermore, let us remark that Theorem 3a implies that whp adding any edge to $G_{p}$, which is not contained in the giant component, must connect an isolated vertex to the giant component, and indeed this suffices to show that the hitting times of minimum degree one and connectivity are whp the same. Theorem 3b then shows that whp the only obstacle to a perfect matching in $G_{p}$ itself is the existence of isolated vertices. Thus, by standard coupling between the random graph process and the bond percolation model, Theorem 2 indeed follows from Theorem 3.
Let us further remark that the proof of Theorem 3a does not rely on the product structure of the graph, and only uses the assumptions that the graph is $d$-regular, has optimal edge-expansion for sets of size polynomial in $d$, and has some mild edge-expansion for larger sets. Only in the proof of Theorem 3b we use the product structure explicitly. Moreover, since the hypercube is bipartite, in order to find a perfect matching in $Q^{t}_{p}$, it suffices to verify Hall’s condition. In our setting, however, the graph $G$ is not necessarily bipartite, which means that in order to show the typical existence of a perfect matching, one needs to consider the Tutte-Berge formula. This requires a much more delicate treatment, which is detailed in the proof outline in the subsequent section. This further requires us to develop tools which hold for wider family of graphs (see Lemmas 3.2, 3.3, and 3.4). These could be of independent interest when treating perfect matchings in the setting of percolation on regular graphs.
@@ 1.2 Proofs outline
Throughout the paper we let $C>1$ be an integer, $(H_{i})_{i=1}^{t}$ be a sequence of $d_{i}$-regular and connected graphs with $1<|V(H_{i})|\leq C$, and let $G=\square_{i=1}^{t}H_{i}$ be their product. We let $G=(V,E)$. We call $t$ the dimension of $G$, and given $u=(u_{1},\ldots,u_{t})\in V$ we call $u_{i}$ the $i$-th coordinate of $u$. Furthermore, we denote by $n:=|V|$ the order of $G$, and by $d\coloneqq d(G)$ the degree of $G$, noting once again that $d=\sum_{i=1}^{t}d_{i}=\Theta_{C}(\ln n)$.
@@
$e(S,S^{C})$ $\geq\frac{|S|\cdot e\cdot\ln\left(\frac{n}{|S|}\right)}{C}\qquad\text{ and }$
$e(S,S^{C})$ $\geq|S|(d-(C-1)\log_{C}|S|).$
@@
The proof follows from a double-exposure argument, similar in spirit to the classical argument of Ajtai, Komlós, and Szemerédi [1]. Let $p$ be as in the statement of Theorem 3. Let $p_{2}\coloneqq\frac{1}{d^{2}}$ and let $p_{1}$ be such that $(1-p_{1})(1-p_{2})=1-p$. Note that $G_{p_{1}}\cup G_{p_{2}}\sim G_{p}$, and $p_{1}\approx p-\frac{1}{d^{2}}$. We first show, using a first-moment calculation, that whp every two vertices $u,v$ which are isolated in $G_{p_{1}}$ are at distance at least two in $G$ (Lemma 2.1). Then, utilising Theorem 1.3 and Lemma 1.4, we show that whp there are no components in $G_{p_{1}}$ whose order is in $\left[2,d^{20}\right]$ (Lemma 2.2). We then turn to show that typically components whose order is at least $d^{20}$ merge after sprinkling with $p_{2}$ — this is fairly standard, and follows the same approach as in [1]. Noting that any vertex which was an isolated vertex in $G_{p_{1}}$, is whp either an isolated vertex after sprinkling with $p_{2}$ or merged into a component which was of size at least $d^{20}$ in $G_{p_{1}}$, we obtain that whp $G_{p}$ has a unique large connected component, and all the other vertices are isolated vertices. Finally, from Markov’s inequality it follows that the total volume of isolated vertices in $G_{p}$ is $o(n)$.
@@
Recall that we seek to show that the giant component of $G_{p}$ has a (nearly-)perfect matching. We assume divisibility by $2$ whenever necessary, noting that minor modifications will allow us to obtain a nearly-perfect matching otherwise. We know from Theorem 3a that whp every vertex of $G$ which is not in the giant component is an isolated vertex in $G_{p}$. Thus, it suffices to show that whp $G_{p}$ has a perfect matching if there are typically no isolated vertices in $G_{p}$. To that end, we will utilise the Tutte-Berge formula [2], which implies that the number of unmatched vertices in a maximum matching in a graph $\Gamma$ is equal to
@@
We thus define Tutte-like obstructions — we say that $U\subseteq V$ is an obstruction (for $G_{p}$) if $|U|\geq 1$, and the number of components of size different than two in $G_{p}[V\setminus U]$ is at least $|U|+1$. We say that the size of the obstruction is $|U|$. Note that if $G_{p}$ has no such obstructions, then the only possible obstructions to a perfect matching are isolated vertices, and whp the giant component has a perfect matching. Furthermore, observe that this definition does not capture all the properties coming from the Tutte-Berge formula – indeed, we treat components with an even number of vertices, except two, the same as components with an odd number of vertices – but this crude, yet much-simplified, outlook will (almost) suffice for our needs.
@@
We say that $U$ is a minimal obstruction, if it is an obstruction with the smallest size of $U$. Note that if a graph has an obstruction, then it must also have a minimal obstruction.
@@
We will use the convention that $u\coloneqq|U|$. In order to obtain a finer description of the components in $G_{p}[V\setminus U]$ we denote the components of size two in $G_{p}[V\setminus U]$ by $W_{1},\ldots,W_{w/2}$, such that their total volume is $w$. Let us partition the vertices in the components that are not of size two into three sets, according to the size of the component in $G_{p}[V\setminus U]$ they belong to. Let $\ell_{1}$ be the number of components of size one in $G_{p}[V\setminus U]$, let $\ell_{2}$ be the number of components whose size is in $3,\frac{n}{d^{C^{3}/p}}]$ in $G_{p}[V\setminus U]$, and let $\ell_{3}=\ell-\ell_{1}-\ell_{2}$ be the number of components in $G_{p}[V\setminus U]$ that contain more than $\frac{n}{d^{C^{3}/p}}$ vertices. We further denote the set of vertices in components of $G_{p}[V\setminus U]$ whose size is one by $V_{1}$, those whose size is in $3,\frac{n}{d^{C^{3}/p}}]$ by $S$, and those whose size is larger than $\frac{n}{d^{C^{3}/p}}$ by $B$. We write $s:=|S|$ and $b:=|B|$, and note that $\ell_{1}=|V_{1}|$.
We remark that the choice of the values for defining $S$ and $B$ is not (completely) arbitrary, and we will see these come into play throughout the proof.
@@
Figure 1: Illustration of an obstruction, with the sets $U$, $V_{1}$, $W$, $S$ and $B$. Note that the only edges in $G_{p}$, which are not induced by $U$ or by components from $W,S$ or $B$, are in $E(U,V_{1}\cup W\cup S\cup B)$.
Using 1.3 one can deduce a lower bound on the number of edges leaving components in $G_{p}[V\setminus U]$. Note that for $U$ to be an obstruction, for any component $K$ of $G_{p}[V\setminus U]$ the edges $E_{G_{p}}(K,K^{C})\subseteq E(K,U)$. Since $G$ is $d$-regular, there are at most $du$ edges touching $U$. So in order to have an obstruction in $G_{p}$, many of the edges leaving components in $G_{p}[V\setminus U]$ cannot be present in $G_{p}$.
We will then treat minimal obstructions differently, depending on their size (that is, on the size of $u$). For $u\leq d$, we will have a small caveat. We call an obstruction $U$ with $\ell=u+1,\ell_{1}=u$ and $\ell_{2}+\ell_{3}=1$ a trivial obstruction. Note that any trivial obstruction is not a real obstruction to a perfect matching in the Tutte-Berge sense as the single component in $S\cup B$ is actually of even size ($n-w-\ell_{1}-u$ is even since $n,w$ are even and $\ell_{1}=u$). Thus, we may assume that any obstruction we consider is not a trivial obstruction. Using this observation, we will be able to show that whp there are no minimal non-trivial obstructions when $u\leq d$ (Lemma 3.10).
@@
Turning to obstructions of more substantial size, we note that an obstruction is uniquely determined by the choice of $U$, and that $U$ is not necessarily connected, and thus the number of choices could be $\binom{n}{u}$. As is often the case, a key ingredient will be to efficiently enumerate the number of possible obstructions, and $\binom{n}{u}$ will often be an inefficient bound. We thus prove several claims, which could be of independent interest in arguing about perfect matchings under percolation for general graphs. In Lemma 3.3, we show that a minimal obstruction of size $u$ is determined by the choice of $W\cup S\cup B$. We stress that this result holds for any graph $G$, without requirements on its degree or structure. As these sets contain connected components, by Lemma 1.4 there are at most (roughly) $n^{w/2+\ell_{2}+\ell_{3}}(ed)^{s+w+b}$ choices for a minimal obstruction of size $u$ (see Lemma 3.4 for a precise statement). Since $d=\Theta_{C}(\ln n)$, for ‘large’ $u$ this will be a much more efficient bound than $\binom{n}{u}$. Furthermore, if $U$ is a minimal obstruction, then every vertex $v\in U$ has neighbours in at least three components of $G_{p}[V\setminus U]$ (see Lemma 3.2). For certain values of $u,\ell_{1},\ell_{2},\ell_{3}$, this implies that typically an obstruction $U$ is contained in a well-connected set of the same order (see Lemma 3.14 for a precise statement). This allows for an even more accurate enumeration. This more effecient counting approach will allow us to treat obstructions up to size $u=n/poly(d)$.